OpenStudy (anonymous):

Where does the graph of y = 2x2 + 7x + 3 cross the x-axis?

OpenStudy (saifoo.khan):

Use the discriminant to find out. P.S. discriminant => b^2 - 4ac

OpenStudy (zzr0ck3r):

set it equal to 0 2x^2 + 7x + 3 2x^2+x+6x+3 (I used the ac method 2(3) = 6 1(6) = 6 and 6+1=7 factor x(2x+1)+3(2x+1)=o (2x+1)(x+3)=0 2x+1 = 0 x = -1/2 x+3 = 0 x = -3

OpenStudy (zzr0ck3r):

what are you doing in class? are you doing the AC method, the (-b+-sqrt(b^2-4ac))/(2a), or completing the square?

OpenStudy (anonymous):

a (-1/2,0) and (-3,0) b (1/2,0) and (3,0) c (-3/2,0) and (-1,0) d (3/2,0) and (1,0)

OpenStudy (zzr0ck3r):

we know the y is going to be 0 because thats how we find x intercepts, so if one x = -1/2 and one x = -3 and the y is going to be 0. which one is it?